# On self-dualities for scalar $\phi^{4}$ theory

### v1

@@ III Chang duality in $d=2$
$\tilde{\Omega}^{(2)}_{R1}=\frac{\tilde{M}^{2}}{8\pi}\left(\frac{1}{\varepsilon}+\frac{\pi\tilde{m}_{B}^{2}}{3\tilde{\lambda}_{B}}+\ln\frac{\bar{\mu}^{2}e^{1}}{\tilde{M}^{2}}\right)+\frac{\tilde{M}^{4}-2\tilde{m}_{B}^{4}}{48\tilde{\lambda}_{B}}\,$, (11)
@@
$\tilde{\Omega}^{(2)}_{R1}=\frac{\tilde{M}^{2}}{8\pi}\ln\frac{\Lambda_{\overline{\rm MS}}^{2}e^{1}}{\tilde{M}^{2}}+\frac{\tilde{M}^{4}-2\tilde{m}_{B}^{4}}{48\tilde{\lambda}_{B}}\,$, (12)
@@ VII.1 Expansion around symmetric phase saddle
$e^{-\lambda_{B}\phi^{4}(x)}=\int\frac{d\zeta}{4\sqrt{\lambda_{B}\pi}}e^{-\frac{\zeta^{2}}{16\lambda_{B}}+i\zeta\phi^{2}(x)}\,$, (33)
@@
Writing $\zeta(x)=\zeta_{0}+\xi(x)$, the zeroth order R-level resummation consists of neglecting fluctuations $\xi(x)$ in the evaluation of the free energy. (This corresponds to the leading large N behavior of an N-component scalar field theory, see [7] for a recent pedagogical review.) The fact that $\zeta_{0}$ is constant makes the R0-level free energy analytically accessible, and one finds

### v2

@@ III Chang duality in $d=2$
$\tilde{\Omega}^{(2)}_{R1}=\frac{\tilde{M}^{2}}{8\pi}\left(\frac{1}{\varepsilon}+\frac{\pi\tilde{m}_{B}^{2}}{3\tilde{\lambda}_{B}}+\ln\frac{\bar{\mu}^{2}e^{1}}{\tilde{M}^{2}}\right)+\frac{\tilde{M}^{4}-2\tilde{m}_{B}^{4}}{96\tilde{\lambda}_{B}}\,$, (11)
@@
$\tilde{\Omega}^{(2)}_{R1}=\frac{\tilde{M}^{2}}{8\pi}\ln\frac{\Lambda_{\overline{\rm MS}}^{2}e^{1}}{\tilde{M}^{2}}+\frac{\tilde{M}^{4}-2\tilde{m}_{B}^{4}}{96\tilde{\lambda}_{B}}\,$, (12)
@@ VII.1 Expansion around symmetric phase saddle
$e^{-\lambda_{B}\phi^{4}(x)}=\int\frac{d\zeta}{4\sqrt{\lambda_{B}\pi}}e^{-\frac{\zeta^{2}}{16\lambda_{B}}+\frac{i\zeta\phi^{2}(x)}{2}}\,$, (33)
@@
Writing $\zeta(x)=\zeta_{0}+\xi(x)$, the zeroth order R-level resummation consists of neglecting fluctuations $\xi(x)$ in the evaluation of the free energy. (This corresponds to the leading large N behavior of an N-component scalar field theory[18], see [7] for a recent pedagogical review.) The fact that $\zeta_{0}$ is constant makes the R0-level free energy analytically accessible, and one finds